What should you pay for a $1,000 10-year 10% Coupon rate bond paid semi-annually, when the required return market rate is 12%?

A. $1,000.00

B.$ 885.30

C.$1,389.73
Answer #1

Option B.$ 885.30

Cash that should be paid Amount Calculation
Present Value of Interest       573.50 =(1,000*10%*6/12)*(1-(1/(1+6%)^20))/6%
Present Value of Maturity       311.80 =1,000/(1+6%)^20
Cash that should be paid       885.30
Interest rate - annual 12%
Interest rate - semi annual 6%
Number of years 10
Number of semi annual periods 20
Formula
Present Value of Interest
PVI = PMT [(1 - (1 / (1 + r)n)) / r]
PVI = Present value of interest to be paid
PMT = Amount of interest payment
r = Interest rate
n = Number of periods over which payments are to be made
Present Value of Maturity
PVM =FV/(1+r)^n
PVM = Present value of maturity amount to be paid
FV = Amount of Maturity payment
r = Interest rate
n = Number of periods over which payments are to be made